OpenStudy (anonymous):

A bulldozer moves 640 lb of dirt in 4/5 h. Complete the unit rate. lb/h

OpenStudy (anonymous):

32

OpenStudy (anonymous):

so what u get+

OpenStudy (agl202):

x=(1)*(640)/(4/5)=640/4*5=160*5=800